OpenStudy (chacha1234):

Please help with this quick 8th grade math question it's due tomorrow please will fan and medal! In the system of equations, which variable would it be easiest to solve for? 2x+4y=26 3x+y=9 A. The easiest to solve for is x in the first equation. B. The easiest to solve for is y in the first equation. C. The easiest to solve for is x in the second equation. D. The easiest to solve for is y in the second equation.

OpenStudy (chacha1234):

@phi @rebeccaxhawaii @freckles @zepdrix @MathMusician

OpenStudy (chacha1234):

Help me out here @zepdrix ?

zepdrix (zepdrix):

It's usually easiest to solve for a variable when it `does not have a coefficient`, a number in front of it. So ya... any ideas on the answer? :)

OpenStudy (chacha1234):

I honestly have no idea but when I solve, I usually solve for x first but it's asking for the easiest and I'm not sure what's the easiest way

OpenStudy (chacha1234):

I think it's b

zepdrix (zepdrix):

Well I just gave you a big hint :O There is only one variable present without a number attached to it.

OpenStudy (chacha1234):

I'm so confused 0=

zepdrix (zepdrix):

really? :\

OpenStudy (chacha1234):

Yep.

zepdrix (zepdrix):

In the first equation... the x has a 2 and the y has a 4.

OpenStudy (chacha1234):

ok..?

zepdrix (zepdrix):

In the second equation... the x has a 3 and the y has no coefficient.

OpenStudy (chacha1234):

So d!

zepdrix (zepdrix):

Yay good job

OpenStudy (chacha1234):

One more please? @zepdrix

OpenStudy (chacha1234):

Which is true regarding the system of equations? 6x+2y=46 3x+y=23 A. The system results in a false statement. B. The system results in an intersection at one point. C. The system results in parallel lines. D. The system results in a true statement because they are the same line.

zepdrix (zepdrix):

If we multiply both sides of the `second equation` by 2, we will have this system, 6x+2y=46 6x+2y=46

zepdrix (zepdrix):

If we combine these equations, maybe subtracting the second equation from the first, what are we left with?

OpenStudy (chacha1234):

? =(

OpenStudy (chacha1234):

All I know is that the equation has infinite solutions.

zepdrix (zepdrix):

Ok well maybe read your options again. When we were able to get to this point, 6x+2y=46 6x+2y=46 One of the options should jump out at you :)

OpenStudy (chacha1234):

No idea..

zepdrix (zepdrix):

They are exactly the same, yes?

OpenStudy (chacha1234):

Yes.

zepdrix (zepdrix):

Hmm, one of your options mentions that...

OpenStudy (chacha1234):

D?